# ACDiT: Interpolating Autoregressive Conditional Modeling and Diffusion Transformer

### v2

@@ 5.3 Analysis
Trade off of block size. Fig. 5 illustrates the trend of trade-off under different sequence lengths and block sizes in image and video generation tasks on ACDiT-B. The FID curve indicates that for image generation, directly increasing the autoregressive length leads to a decline in image quality, since each patch receives less attention information on average. However, we can mitigate this decline by increasing the total sequence length, which means reducing the patch size. For video generation, ACDiT shows more advantages due to the inherent temporal dependence of videos. The FVD curve demonstrates that increasing autoregressive length has minimal effect on the video quality, even with slight improvement. As for efficiency, we test the sampling time for various sequence lengths with a batch size of 4 on an NVIDIA A100 GPU. Fig. 6 shows that as the sequence length increases, particularly beyond 16k, full-sequence attention (AR length of 1) becomes very time-consuming, necessitating the autoregressive generation.
@@
Scaling Performance. We present the scaling performance of ACDiT in Fig. 7(a). For a fair comparison with DiT, we use the same batch size and learning rate as DiT in these training sessions. When increasing the model size, ACDiT shows consistent improvement in image quality across all autoregressive lengths, sharing a similar scaling trend with DiT. Notably, the improvement is more pronounced with longer autoregressive lengths. We hypothesize that this is due to reduced accumulation of errors when scaling the model size.
Ablation Study. We ablate the effectiveness of ROPE-ND positional embedding on image generation with patch size as 2. As shown in Fig 7(b), adding ROPE-ND results in consistent improvements in FID score.
@@
Training dynamics of WSD scheduler. Unlike the constant learning rate used in DiT, we utilize the WSD learning rate scheduler [33]. WSD scheduler maintains a constant learning rate as the main stage of training, while one can diverge from the main branch at any time, potentially based on the compute budget, with a rapidly decaying learning rate. As Fig. 7(c) shows, the FID remains almost converged during the constant learning rate state, while sharply dropping after the learning decays, similar to the loss curve when using the WSD scheduler in LLM training. To the best of our knowledge, we are the first to validate the effectiveness of the WSD scheduler in visual generation.

### v3

@@ 5.2 Analysis
Trade off of block size. Fig. 4 illustrates the trend of trade-off under different sequence lengths and block sizes in image and video generation tasks on ACDiT-B. The FID curve indicates that for image generation, directly increasing the autoregressive length leads to a decline in image quality, since each patch receives less attention information on average. However, we can mitigate this decline by increasing the total sequence length, which means reducing the patch size. For video generation, ACDiT shows more advantages due to the inherent temporal dependence of videos. The FVD curve demonstrates that increasing autoregressive length has minimal effect on the video quality, even with a slight improvement. As for efficiency, we test the sampling time for various sequence lengths with a batch size of 4 on an NVIDIA A100 GPU. Fig. 5 shows that as the sequence length increases, particularly beyond 16k, full-sequence attention (AR length of 1) becomes very time-consuming, necessitating the autoregressive generation. Empirically, blocks spanning 2–4 frames provide a good quality–efficiency trade-off for video generation. Dynamic or content-adaptive block sizing is a promising extension, and we leave it for future work.
Inference Cost Comparison with MAR. We further compare the average inference time of ACDiT and MAR for image generation and video generation in Table 6. Since MAR is not designed or trained for video generation, both models are evaluated under a controlled setting with a fixed sequence length of 4096 tokens to reflect inference cost in a long-horizon setting. Notably, ACDiT is consistently faster than MAR even for standard image generation. This shows that KV-cached block-wise autoregression already improves efficiency for common image-generation workloads. The advantage becomes substantially larger for long sequences. While MAR recomputes full bidirectional attention at each generation step, ACDiT caches key–value pairs of clean past blocks and only performs attention on newly generated blocks, avoiding redundant computation and achieving much lower (10x) latency for long-horizon video generation.
@@
Scaling Performance. We present the scaling performance of ACDiT in Fig. 6(a). For a fair comparison with DiT, we use the same batch size and learning rate as DiT in these training sessions. When increasing the model size, ACDiT shows consistent improvement in image quality across all autoregressive lengths, sharing a similar scaling trend with DiT. Notably, the performance gap between different autoregressive lengths shrinks as model capacity grows. This suggests that larger models can more accurately fit each autoregressive conditional distribution, thereby reducing the accumulation of errors over long autoregressive horizons.
ROPE-ND. We ablate the effectiveness of ROPE-ND on image generation. As shown in Fig 6(b), both removing ROPE-ND and replacing RoPE-ND with standard 1D-RoPE consistently degrade FID.
Effect of clean-past conditioning. We further ablate the effect of clean-past conditioning by replacing clean context latents with noisy ones when modeling autoregressive conditionals. As shown in Fig. 6(b), removing clean-past conditioning leads to consistently worse generation quality. Beyond generation, using clean latents also provides benefits for representation learning, as evidenced by the improved downstream classification performance discussed in Table 4. These results indicate that clean-past conditioning is beneficial for both autoregressive generation and learning informative visual representations.
@@
Training dynamics of WSD scheduler. We utilize the WSD learning rate scheduler (Hu et al., 2024). WSD scheduler uses a constant learning rate during main training, while allowing divergence at any point with a rapidly decaying learning rate based on compute budget. As Fig. 6(c) shows, the FID remains almost converged during the constant learning rate state, while sharply dropping after the learning decays, similar to the loss curve when using the WSD scheduler in LLM training. To the best of our knowledge, we are the first to validate the effectiveness of the WSD scheduler in visual generation.
@@ 6 Conclusion
Appendix A Detailed Discussion about Block Size
In this section, we provide a detailed analysis of the computational cost of ACDiT in terms of FLOPS, comparing it to standard full-sequence diffusion in transformer layers. Let $L$ denote the sequence length, $h$ the hidden size, $n$ the number of attention heads (assumed $n\ll h$), $\theta$ the number of parameters in a single transformer layer, and $B$ the block size used in ACDiT.
Full-Sequence FLOPS.
Each transformer layer has two major computational components:
  • Linear projections and feed-forward layers: $2L\theta$
  • Q–K attention dot-products: $4(h+n)L^{2}\approx 4hL^{2}$
Hence, the total FLOPS per layer is:
$F_{\text{full}}=2L\theta+4hL^{2}$
ACDiT with KV-Cache.
In ACDiT, attention is computed block-wise using KV-Cache. The $i$-th block (of size $B$) attends to (i1)B(i-1)B cached tokens, requiring $iB^{2}$ Q–K dot-products. Summing over all $\tfrac{L}{B}$ blocks:
$\sum_{i=1}^{L/B}iB^{2}=B^{2}\cdot\frac{L}{B}\cdot\left(\frac{L}{B}+1\right)\bigg/2=\frac{L^{2}+LB}{2}$
So the attention FLOPS becomes:
$\text{Attention}_{\textsc{ACDiT}}=4h\cdot\frac{L^{2}+LB}{2}=2hL^{2}+2hLB$
Including projection and FFN cost, the total FLOPS under ACDiT is:
$F_{\textsc{ACDiT}}=2L\theta+2hL^{2}+2hLB$
FLOPS Savings.
The absolute FLOPS saved per layer is:
$F_{\text{saved}}=F_{\text{full}}-F_{\textsc{ACDiT}}=2hL^{2}(1-\tfrac{B}{L})$
The relative savings are:
$\frac{F_{\text{saved}}}{F_{\text{full}}}=\frac{2hL^{2}(1-\frac{B}{L})}{2L\theta+4hL^{2}}=\frac{1-\tfrac{B}{L}}{2+\tfrac{\theta}{hL}}$
When $L\gg h$, the savings approaches $\tfrac{1}{2}(1-\tfrac{B}{L})$, implying up to 50% FLOPS reduction when $B\ll L$.
In practice, it is not always beneficial for small $B$. Setting $B$ to an excessively small value may not fully leverage the iterative modification inherent in the diffusion process, potentially compromising generation quality. Furthermore, given the parallel computing nature of computational kernels, a very small $B$ may not yield speed improvements, a phenomenon analogous to the rationale behind speculative decoding (Leviathan et al., 2023). Conversely, setting $B$ to a very large value diminishes efficiency both in terms of attention calculation. It also fails to capitalize on the strengths of auto-regressive generation. Indeed, when $B$ is set equal to $L$, ACDiT reverts to the original DiT model.
Appendix B Details of Model Architecture and Implementation
ACDiT mainly inherits the main architecture of DiT. We replace the original bidirectional attention with the proposed SCAM attention pattern to process the clean and noisy latent. Since we want to keep the architecture as simple and unified as possible, we use linear layers instead of convolution in the input layer and final layer. Besides, we replace the position embedding and Layer Normalization with RoPE (Su et al., 2024) and RMSNorm (Root Mean Square Layer Normalization) (Touvron et al., 2023a), respectively. We find that QK-norm is important to stabilizing the video generation training, thus we use QK-norms in all experiments. The additional conditional information, including timesteps and labels, is injected into the model with AdaLN-Zero only on the noise part. For both image and video generation, we follow DiT and leverage the pre-trained image VAE (Kingma, 2013) from Stable Diffusion (Rombach et al., 2022), whose downsample factor is 8. For image generation under B block size, we group square latent representation patches with $\sqrt{B}\times\sqrt{B}$ shape as a block. We do not further train the tokenizer in the target dataset since it’s out of the scope of our work, despite that further training might yield better results.
In the image generation task, we set the patch size as 1 and the autoregressive unit block size as $256=16\times 16$. Therefore, for a $256\times 256\times 3$ image in $32\times 32\times 4$ latent shape, the total sequence length and autoregressive length are 1024 and 4, respectively. We explore 4 different model sizes, as shown in Table 6. ACDiT-B is used for design verification and analysis. ACDiT is trained on ImageNet for 1.2M iterations with a batch size of 1024. We use the AdamW optimizer (Loshchilov, 2017) and WSD (Warmup Steady Decay) learning rate scheduling (Hu et al., 2024) with the peak learning rate 3e-4 and no weight decay. The learning rate begins to decay in the last 15% training iterations. Following the common training recipe of generative models, we keep an exponential moving average (EMA) of the ACDiT weights during training using a decay rate of 0.9999. We sample images with DPM-Solver (Lu et al., 2022) for 25 steps within each block and use classifier-free guidance (Ho and Salimans, 2022) with a guidance scale of 1.5. In video generation, we sample 16 frames from each video and set the patch size as 2 and the block size as $1024=256\times 4$. For a $16\times 256\times 256\times 3$ video in $16\times 32\times 32\times 4$ latent shape, the sequence length of each frame is 256 and the total sequence length is 4096, with 4 frames grouped into one block. We train ACDiT on UCF-101 for 400K iterations with a batch size of 96. The classifier-free guidance scale is 2.5. Other training configs are the same as image training. All models are implemented with PyTorch (Paszke et al., 2019). Specifically, we use FlexAttention [^3] https://pytorch.org/blog/flexattention to implement the SCAM for both customization and efficiency. We use 256 and 96 NVIDIA H100 GPUs to train the image and video generation model, respectively.
Appendix C ACDiT PyTorch Code Example
In this section, we provide key implementation details of ACDiT, including (1) how clean and noisy blocks are arranged in the sequence, (2) how the Skip-Causal Attention Mask (SCAM) is constructed, and (3) how SCAM integrates with the attention module and KV-cache for efficient autoregressive inference.
For $N$ autoregressive blocks, we build a sequence of length $2N$ containing all clean blocks followed by all noisy blocks:
$\texttt{idx\_clean}(i)=i,\quad\texttt{idx\_noisy}(i)=N+i,\quad i=0,\dots,N-1.$
The following code shows how to construct SCAM, where noisy blocks are only allowed to attend to all previous clean blocks and themselves, while clean blocks attend to all preceding clean blocks. This attention mask enforces clean-past conditioning and enables KV-cache reuse across blocks.
1def build_attention_mask(self, T, B, device):
2 size = T * B
3 m_noise_noise = torch.zeros(size, size)
4 for i in range(T):
5 start_idx = i * B
6 end_idx = start_idx + B
7 m_noise_noise[start_idx:end_idx, start_idx:end_idx] = torch.ones(B, B)
8 m_noise_clean = torch.zeros(size, size)
9 for i in range(T):
10 for j in range(i + 1, T):
11 start_col = i * B
12 end_col = start_col + B
13 start_row = j * B
14 end_row = start_row + B
15 m_noise_clean[start_row:end_row, start_col:end_col] = 1
16 m_clean_noise = torch.zeros(size, size)
17 m_clean_clean = torch.zeros(size, size)
18 for i in range(T):
19 start_idx = i * B
20 end_idx = start_idx + B
21 m_clean_clean[start_idx:end_idx, :end_idx] = 1
22
23 attn_mask = torch.zeros(2 * size, 2 * size)
24 attn_mask[:size, :size] = m_clean_clean
25 attn_mask[:size, size:] = m_clean_noise
26 attn_mask[size:, :size] = m_noise_clean
27 attn_mask[size:, size:] = m_noise_noise
28 return attn_mask.bool().to(device)
If using PyTorch FlexAttention, the same causal dependency can be implemented through its index-based mask generation. Below is the corresponding function:
1def skip_causal_attn_mask_mod_gen(b, h, q_idx, kv_idx, block_size, len1):
2 q_idx = q_idx // block_size
3 kv_idx = kv_idx // block_size
4 mask = torch.where(((q_idx < len1) & (kv_idx < len1) & (q_idx >= kv_idx))| ((q_idx >= len1) & (kv_idx < len1) & ((q_idx - len1) > kv_idx)) | (q_idx == kv_idx) , True, False)
5 return mask
Once the mask is constructed, it is passed to the attention module. KV-cache is applied only to clean blocks, allowing us to cache context features while discarding the noisy-block activations after each diffusion step.
1class SkipCausalAttention(Attention):
2 def forward(self, x, position_ids=None, attention_mask=None, block_size=None, cache=False):
3 B, N, C = x.shape
4 qkv = self.qkv(x).reshape(B, N, 3, self.num_heads, self.head_dim).permute(2, 0, 3, 1, 4)
5 q, k, v = qkv.unbind(0)
6 q, k = self.q_norm(q), self.k_norm(k)
7 if self.rope is not None:
8 q, k = self.rope(q, k, position_ids)
9 if self.caching:
10 if cache:
11 if self.cached_k is None:
12 self.cached_k = k[:, :, :block_size, :]
13 self.cached_v = v[:, :, :block_size, :]
14 self.cached_x = x
15 else:
16 self.cached_k = torch.cat((self.cached_k, k[:, :, :block_size, :]), dim=2)
17 self.cached_v = torch.cat((self.cached_v, v[:, :, :block_size, :]), dim=2)
18 if self.cached_k is not None:
19 k = torch.cat((self.cached_k, k[:, :, -block_size:, :]), dim=2)
20 v = torch.cat((self.cached_v, v[:, :, -block_size:, :]), dim=2)
21 if not USE_FLEX_ATTENTION:
22 x = torch.nn.functional.scaled_dot_product_attention(
23 q, k, v, attn_mask=attention_mask, dropout_p=self.attn_drop.p
24 )
25 else:
26 x = flex_attention(q, k, v, block_mask=attention_mask)
27 x = x.transpose(1, 2).reshape(B, N, C)
28 x = self.proj(x)
29 x = self.proj_drop(x)
30 return x
ACDiT uses a standard Transformer block design as in DiT, with one key modification: only the noisy latent tokens interact with the conditional information (diffusion timestep, class label). Clean latents remain unconditioned to preserve their deterministic role as the cached context.
1class ACDiTBlock($n$.Module):
2 def forward(self, x, c, attention_mask=None, cond_length=0, block_size=None, cache=False, position_ids=None):
3 N, T, _, C = c.shape
4 N, TB, C = x[:, cond_length:].shape
5 B = TB // T
6
7 ada_c_list = self.adaLN_modulation(c).chunk(6, dim=-1)
8 (shift_msa, scale_msa, gate_msa, shift_mlp, scale_mlp, gate_mlp) = [ada_c.repeat(1, 1, B, 1).reshape(N, TB, C)for ada_c in ada_c_list]
9 norm_x1 = self.norm1(x.to(torch.float32)).to(dtype)
10 attn_input_x = torch.cat((norm_x1[:, :cond_length], modulate(norm_x1[:, cond_length:], shift_msa, scale_msa))dim=1)
11 attn_output_x = self.attn(attn_input_x, attention_mask=attention_mask, block_size=block_size, cache=cache, position_ids=position_ids)
12 x = x + torch.cat((attn_output_x[:, :cond_length], gate_msa * attn_output_x[:, cond_length:]), dim=1)
13
14 norm_x2 = self.norm2(x.to(torch.float32)).to(dtype)
15 gate_input_x = torch.cat((norm_x2[:, :cond_length], modulate(norm_x2[:, cond_length:], shift_mlp, scale_mlp)), dim=1)
16 gate_output_x = self.mlp(gate_input_x)
17 x = x + torch.cat((gate_output_x[:, :cond_length], gate_mlp * gate_output_x[:, cond_length:]), dim=1)
18 return x
Appendix D Broader Impact
This work advances the efficiency and scalability of generative models for images, videos, and text by combining autoregressive modeling with diffusion. While such models enable positive applications in creative content generation, simulation, and representation learning, they may also be misused to generate misleading or harmful visual or textual content, including deepfakes. As with other general-purpose generative models, responsible deployment requires appropriate safeguards, dataset curation, and usage policies. We encourage future work to explore detection, watermarking, and alignment techniques alongside improvements in generative quality and efficiency.
Appendix E Qualitative Results of ACDiT
We show the qualitative results of ACDiT in Figure 7.
Figure 7: Sample images from ACDiT-H on ImageNet $\times256$ and sample videos from ACDiT-XL trained on UCF-101.